A firm has the production function f(x1, x2) = x11x0.502. The isoquant on which output is 305/10 has the equation

a.

x2 = 30x-0.501.

b.

x1/x2 = 2.

c.

x1 = 0.50x-0.502.

d.

x2 = 30x-21.

e.

x2 = 30x21.

step by step, plese
Answer #1

Given that output is determined by the equation

f(x1, x2) = x1^1 x2^0.5

Q =  x1 * x2^0.5

30^5/10 =  x1 * x2^0.5

30^0.5 =  x1 * x2^0.5

x2^0.5 = 30^0.5 / x1

x2 = (30^0.5 / x1)^2

x2 = 30^(0.5*2) / x1^2

x2 = 30/x1^2

x2 = 30x1^-2

Hence the correct option is d).